# Material Point Methods on Unstructured Tessellations: A Stable Kernel Approach With Continuous Gradient Reconstruction

### v1

@@ 4 Conclusion, Limitations, and Future Work
This study has extended the Moving Least Square Material Point Method (UMLS-MPM) to encompass general tessellations within both 2D and 3D meshes. This advance has been achieved through the multiplication of a diminishing function to the MLS sample weights. Analytically proved, the approach ensures continuous kernel reconstruction and provides a sound foundation for high-order MPM on any unstructured mesh types. Several numerical experiments in both 2D and 3D domains have shown the method’s effectiveness in high-order convergence and the elimination of cell-crossing errors.
@@
However, the proposed method still has limitations. To ensure the MLS solution does not degenerate, there are quality requirements on the surrounding vertices of the particle, which pose challenges, especially in meshes with sharp resolution transitions or poor quality. Another trigger for degeneracy is the lack of surrounding samples on one side; for example, when the particle is near the domain boundary. While this issue was mitigated in our last experiment by drawing an extra layer of vertices, their automatic creation via an algorithm would be an improvement. Additionally, the selection of sample weights also poses a challenge. We used a B-spline function with a fixed unit support length equaling the maximum element size to ensure that the first ring of neighbors is reserved even in the coarsest area. However, this strategy leads to more uniform sample weights in finer regions, blurring the kernel. A potential solution can be to incorporate a sizing field within the mesh to dynamically adjust the sample weight function.
@@
From an application standpoint, exploring the integration between the Material Point Method (MPM) and gas or fluid simulations via the Finite Volume Method (FVM) presents significant potential [21, 53]. Furthermore, recent advances have seen learning MPM or other mesh-based simulations using graph neural networks, accelerating the inferences [54, 55, 56]. Our kernel construction suggests a potentially novel learning paradigm for MPM on unstructured meshes, which is similar to embedding both kernel and mesh information into the network’s channel [57, 58].
Appendix A Proofs for the Continuous Reconstructions
A.1 Proofs
@@
For conciseness, we drop the subscripts ${}_{vp}$ in the following proofs. We start by assuming there exists a smooth, diminishing function $\eta$ for the nodes added or removed from the set of nearby nodes ${\mathcal{N}}^{1}$ when a particle crosses the boundary of a cell. Under this assumption, we can prove that our kernel value and gradient estimation is continuous across the boundary. We present the proof in 2D when a particle crosses an edge; the extension to 3D and other crossing cases is straightforward. Then, we prove that the proposed function in (23) is locally diminishing for ${\mathcal{N}}_{a,r}^{1}$ when the particle crosses the edge.
Proposition 1.
@@
Our kernel value and gradient estimation is continuous across cell boundaries.
@@
Let ${\mathcal{N}}_{o,n}^{1}$ be the sets of nearby nodes before/after the particle $p$ crosses the common edge between the old/new cells ${\mathcal{N}}_{o,n}^{0}$. Here, the subscripts ${o,n}$ denote the old/new cell, respectively, and the superscripts ${0,1}$ indicate the ring-0/1 neighbors of the cell, respectively. Let ${\bm{x}}_{p}^{o,n}$ be the position of particle $p$ before/after the crossing and $||{\bm{x}}_{p}^{n}-{\bm{x}}_{p}^{o}||={\mathcal{O}}(\epsilon)$. Define the common node set ${\mathcal{N}}_{c}^{1}={\mathcal{N}}_{o}^{1}\cap{\mathcal{N}}_{n}^{1}$, the added node set ${\mathcal{N}}_{a}^{1}={\mathcal{N}}_{n}^{1}\setminus{\mathcal{N}}_{c}^{1}$, and the removed node set ${\mathcal{N}}_{r}^{1}={\mathcal{N}}_{o}^{1}\setminus{\mathcal{N}}_{c}^{1}$. Since $\eta$ is locally diminishing for $v\in{\mathcal{N}}_{a,r}^{1}$, we have a positive value $K_{1}$ such that $\eta={\mathcal{O}}(K_{1}\epsilon)={\mathcal{O}}(\epsilon)$. The pertubation for the assembled matrix ${\bm{M}}$ before/after the particle $p$ crosses an edge is
@@
Similarly, for the perturbation in the assembled vector ${\bm{Q}}^{T}{\bm{D}}{\bm{u}}$ before/after the particle crossing is
$||\delta\left({\bm{Q}}^{T}{\bm{D}}{\bm{u}}\right)||$ $=||\sum_{v\in{\mathcal{N}}_{c}^{1}}\delta(\eta du{\bm{P}})+\sum_{v\in{\mathcal{N}}_{a}^{1}}\eta du{\bm{P}}-\sum_{v\in{\mathcal{N}}_{r}^{1}}\eta du{\bm{P}}||$ (29)
@@
Furthermore, we can establish the following bound for the assembled vector ${\bm{Q}}^{T}{\bm{D}}{\bm{u}}$:
$||{\bm{Q}}^{T}{\bm{D}}{\bm{u}}||$ $=||\sum_{v\in{\mathcal{N}}^{1}}\eta du{\bm{P}}||$ (30)
@@
Finally, the perturbation for $\left[\hat{u}_{p},\nabla\hat{u}_{p}^{T}\right]^{T}$ from (22) is
$\left[\hat{u}_{p},\nabla\hat{u}_{p}^{T}\right]^{T}$ $=||\delta({\bm{M}}^{-1}{\bm{Q}}^{T}{\bm{D}}{\bm{u}})||$ (31)
$=||\delta{\bm{M}}^{-1}{\bm{Q}}^{T}{\bm{D}}{\bm{u}}+{\bm{M}}^{-1}\delta\left({\bm{Q}}^{T}{\bm{D}}{\bm{u}}\right)||$
$\leq||\delta{\bm{M}}^{-1}{\bm{Q}}^{T}{\bm{D}}{\bm{u}}||+||{\bm{M}}^{-1}\delta\left({\bm{Q}}^{T}{\bm{D}}{\bm{u}}\right)||$
$\leq||\delta{\bm{M}}^{-1}||\cdot||{\bm{Q}}^{T}{\bm{D}}{\bm{u}}||+||{\bm{M}}^{-1}||\cdot||\delta\left({\bm{Q}}^{T}{\bm{D}}{\bm{u}}\right)||$
@@
In the incomplete singular value decomposition of ${\bm{M}}$, the singular values will always be non-negative. And if the surrounding nodes are not degenerate, the minimum singular value $\sigma({\bm{M}})_{\min}$ will always be positive and the condition number of ${\bm{M}}$ is bounded. Therefore, as long as the mesh is of reasonably good quality, both the function value and gradient estimation is ${\mathcal{C}}^{0}$ across the boundary. ∎
Proposition 2.
The function $\eta$ in (23) is locally diminishing for $\forall v\in{\mathcal{N}}_{a,r}^{1}$.
@@
Formally, we need to prove that for any $v\in{\mathcal{N}}_{a,r}^{1}$, when ${\bm{x}}_{p}$ is crossing the edge of a triangle and $||{\bm{x}}_{p}^{n}-{\bm{x}}_{p}^{o}||={\mathcal{O}}(\epsilon)$, the smoothing function $\eta={\mathcal{O}}(\epsilon)$.
Denote the edge that the particle is crossing as $e$ and the portion of $||{\bm{x}}_{p}^{n}-{\bm{x}}_{p}^{o}||$ in the new/old cell as $L^{n,o}$. Trivially,
@@
Then, let the far-away node not on the edge but in the new/old cell be $v_{far}$ (i.e., $v_{far}^{n,o}\notin e\,\land\,v_{far}^{n,o}\in{\mathcal{N}}_{o,n}^{0}$) and the height from a node $v$ to an edge $e$ be $H(v,e)$. Since the height is orthogonal to the edge, we have $H({\bm{x}}_{p}^{n,o},e)\leq L^{n,o}={\mathcal{O}}(\epsilon)$. Consider the barycentric coordinate contributed by the far-away node, in the new/old cell respectively, for ${\bm{x}}_{p}$:
$B^{n,o}_{v_{far}}$ $=\frac{H({\bm{x}}_{p}^{n,o},e)\cdot||e||}{H(v_{far}^{n,o},e)\cdot||e||}$ (33)
$=\frac{H({\bm{x}}_{p}^{n,o},e)}{H(v_{far}^{n,o},e)}$
$={\mathcal{O}}(\frac{\epsilon}{H(v_{far}^{n,o},e)})$
@@
Figure 23: A visual representation of the notations used to prove that $\eta$ diminishes locally, as described in (33), for every vertex $v$ within the first ring of neighbors, ${\mathcal{N}}^{1}_{a,r}$. The dashed line denotes the perpendicular height from a given position to the shared edge. Dotted lines are drawn to construct a triangle between the point ${\bm{x}}_{p}$ and the edge, facilitating the computation of the barycentric coordinates.
Finally, if a node is added/removed during the particle crossing (i.e., $v\in{\mathcal{N}}_{a,r}^{1}$), this means that $v$ is only connected to the far-away nodes $v_{far}^{n,o}$ but not to the edge $e$; i.e., ${A}_{v,v_{far}^{n,o}}=1,\forall v\in{\mathcal{N}}^{1}_{a,r}$, otherwise ${A}_{v,n}=0,\forall n\in e\land\forall v\in{\mathcal{N}}^{1}_{a,r}$. Hence,
$\eta$ $=\sum_{n\in{\mathcal{N}}^{0}}B_{n}{A}_{v,n}$ (34)
$=B^{n,o}_{v_{far}}{A}_{v,v_{far}^{n,o}}+\sum_{n\in e}B_{n}{A}_{v,n}$
$=B^{n,o}_{v_{far}}\cdot 1+\sum_{n\in e}B_{n}\cdot 0$
$=B^{n,o}_{v_{far}}$
$={\mathcal{O}}(\epsilon),\quad\forall v\in{\mathcal{N}}^{1}_{a,r}.$
This concludes the proof. ∎
Appendix B Settings and Analytical Solutions for the Verification Experiments of the Continuous Reconstructions
@@
Appendix C Conservation of the Linear and Affine Momentum When Combined With Affine Particle-in-Cell
@@
$\sum_{i}w_{ip}^{n}$ $=1,$
$\sum_{i}w_{ip}^{n}{\bm{x}}_{i}^{n}$ $={\bm{x}}_{p}^{n},$
$\sum_{i}w_{ip}^{n}({\bm{x}}_{i}^{n}-{\bm{x}}_{p}^{n})$ $=0,$
then the system’s total linear and angular momentum will be conserved when combined with the Affine Particle in Cell (APIC) scheme [59, 60]. A simple introduction to APIC is given here for the sake of completeness, while the detailed proof can found in the supplementary document of the original APIC paper [59].
In APIC, mass $m_{p}$, position ${\bm{x}}_{p}$, velocity ${\bm{v}}_{p}$, and an affine matrix ${\bm{B}}_{p}=\sum_{i}w_{ip}v_{i}({\bm{x}}_{i}-{\bm{x}}_{p})^{T}$ are stored and tracked on particles. Then,
@@
$\begin{split}m_{i}^{n}&=\sum_{p}w_{ip}^{n}m_{p}\\ {\bm{D}}_{p}^{n}&=\sum_{i}w_{ip}^{n}({\bm{x}}_{i}^{n}-{\bm{x}}_{p}^{n})({\bm{x}}_{i}^{n}-{\bm{x}}_{p}^{n})^{T}\\ m_{i}^{n}{\bm{v}}_{i}^{n}&=\sum_{p}w_{ip}^{n}m_{p}({\bm{v}}_{p}^{n}+{\bm{B}}_{p}^{n}({\bm{D}}_{p}^{n})^{-1}({\bm{x}}_{i}^{n}-{\bm{x}}_{p}^{n}))\end{split}$ (37)
@@
$\begin{split}{\bm{v}}_{p}^{n+1}&=\sum_{i}w_{ip}^{n}\tilde{v}_{i}^{n+1}\\ {\bm{B}}_{p}^{n+1}&=\sum_{i}w_{ip}^{n}\tilde{v}_{i}^{n+1}({\bm{x}}_{i}^{n}-{\bm{x}}_{p}^{n})^{T},\end{split}$ (38)

### v2

@@ 4 Conclusion, Limitations, and Future Work
This study has extended the Moving Least Square Material Point Method to encompass general tessellations within both 2D and 3D meshes. This advance has been achieved through the multiplication of a diminishing function to the MLS sample weights. Analytically proved, the approach ensures continuous kernel reconstruction and provides a sound foundation for MPM on any unstructured mesh types. Several numerical experiments in both 2D and 3D domains have demonstrated the method’s effectiveness in achieving high-order convergence and eliminating cell-crossing errors. However, the proposed method still has limitations.
@@
To ensure the MLS solution does not degenerate, there are quality requirements for the surrounding nodes of the particle, which pose challenges, especially in meshes with sharp resolution transitions or poor quality. Additionally, related to varying mesh resolution is the sample weights: we used a B-spline function with a fixed support radius equaling the maximum edge size to ensure that the first ring of neighbors is reserved in the coarsest area. However, this strategy leads to nearly uniform sample weights in finer regions, blurring the kernel. A potential solution could be to incorporate a sizing field within the mesh to dynamically adjust the sample weight function.
Although UMLS-MPM has advantages in conforming to irregular geometry, this also brings potential issues and leaves room for further improvements. For example, the lack of surrounding samples on one side when the particle is inside the boundary cell can lead to kernel degeneration. While this issue was mitigated in our experiments by drawing an extra layer of cells, an automatic and algorithmic approach is more appealing. Combining UMLS-MPM with regular MPM is also promising, as it leverages the advantage of UMLS-MPM in conforming to boundary shapes, as well as the robustness and efficiency of regular MPM for interior domains. For the transition between interior nodes and boundary nodes, methods similar to immersed FEM [65, 46, 43] could be effective, as they project information from regular background grids to irregular meshes.
From a practical standpoint, we need not a higher-order polynomial basis or higher rings of neighbors as long as the current schemes, which have the lowest cost as discussed in Section 2, provides sufficient accuracy. However, from a theoretical perspective, these two prososals are still interesting. Several challenges can be noted: 1. Changing to higher rings of neighbors will require a different solution for the diminishing function (32) to accommodate the new discrete change of active sets. 2. The higher-order polynomial basis will result in a different MLS solution compared to the linear solution (31). Additionally, since the propositions in Appendix B are based on the linear solution, it remains an open question whether the higher-order solution has a continuous kernel and gradient, even with a diminishing function. 3. The higher-order polynomial basis will necessitate a more delicate selection of quadrature point locations. Unlike higher-order FEM simulations, where the quadrature points are usually well-designed and fixed [32], in MPM, these points may need to be adjusted with the particle locations, requiring extra design and attention.
@@
From an application standpoint, exploring the integration between the Material Point Method (MPM) and gas or fluid simulations via the Finite Volume Method (FVM) presents significant potential [7, 3]. For simulating granular materials in realistic and irregular container shapes, combining UMLS-MPM with the Discrete Element Method (DEM) [13] is also an interesting direction. Moreover, object contact detection and handling [42] is crucial to prevent artificial penetration and sticking commonly seen in MPM [36]. In the field of scientific machine learning, recent advances have demonstrated learning MPM or other mesh-based simulations using graph neural networks, accelerating the inferences [49, 41, 8]. Our kernel construction suggests a potentially novel learning paradigm for MPM on unstructured meshes, similar to embedding both kernel and mesh information into the network’s channel [20, 44].
Appendix A Pipeline for Rapid Cell Search
Fast determination of which cell contains a specific particle is crucial for the efficiency of UMLS-MPM. We propose a hash grid-based method to accelerate this process. The pipeline consists of two steps: (1) building the hash grid connectivity table, a dictionary mapping the hash grid as a key to all of its touching cell indices, as detailed in Algorithm 2, and (2) the online search for determining which cell contains a given particle, as described in Algorithm 3.
Algorithm 2 Build Hash to Adjacent Cell
1:Input: Mesh node positions: pos, mesh cells: cell, hash grid size: dx
2:Output: Hash to adjacent cell connectivity: hash2cell
3:Calculate the bounding box of the whole mesh
[…]
6:Initialize hash2cell as an empty dictionary
[…]
8:For each cell, find its bounding box, determine the spatial hash grids it touches, and append the cell to those hash grids
9:for cell_idx in range(len(cell)) do
[…]
16:     for c in candidates do
17:                           hash2cell[c].append(cell_idx)
18:     end for
19:end for
Algorithm 3 Find Containing Cell
1:Input: Mesh node positions: pos, mesh cells: cell, Mesh minimum position: min_pos, hash grid size: dx, Hash to adjacent cell connectivity: hash2cell, particle position: xp
2:Output: Containing cell index or None
3:Get the hash spatial index for the particle
4:          hash_idx $\leftarrow$ floor((xp-min_pos) / dx)
5:Get all adjacent cells to this hash grid
6:          candidates $\leftarrow$ hash2cell[hash_idx]
7:For each candidate cell, check if the particle is inside
8:for c in candidates do
9:     e $\leftarrow$ cell[c]
10:     vs $\leftarrow$ pos[e]
11:     # get the bounding box of the cell
12:     e_min $\leftarrow$ min(vs, axis=0)
13:     e_max $\leftarrow$ max(vs, axis=0)
14:     # quick filter using bounding box
15:     if all(xp >= e_min) and all(xp <= e_max) then
16:         # check barycentric coordinates
17:         bc $\leftarrow$ barycentric_coord(vs, xp)
18:         if all(bc >= 0) then
19:              return c
20:         end if
21:     end if
22:end for
23:# If no candidate cells contain xp
24:return None
Appendix B Proofs for the Continuous Reconstructions
@@
For conciseness, we drop the subscripts ${p}$ in the following proofs. We start by assuming there exists a smooth, locally diminishing function $\eta$ for the nodes added or removed from the set of nearby nodes ${\mathcal{N}}^{1}$ when a particle crosses the boundary of a cell. Under this assumption, we can prove that our kernel value and gradient estimation is continuous across the boundary. We present the proof in 2D when a particle crosses an edge; the extension to 3D and other crossing cases is straightforward. Finally, we prove that (32) satisfies the forementioned assumption.
@@
Proposition. Our kernel value and gradient estimation is continuous across cell boundaries.
@@
Let ${\mathcal{N}}_{o,n}^{1}$ be the sets of nearby nodes before/after the particle $p$ crosses the common edge between the old/new cells ${\mathcal{N}}_{o,n}^{0}$. Here, the subscripts ${o,n}$ denote the old/new cell, respectively, and the superscripts ${0,1}$ indicate the ring-0/1 neighbors of the cell, respectively. Let ${\bm{x}}^{o,n}$ be the position of particle $p$ before/after the crossing and $||{\bm{x}}^{n}-{\bm{x}}^{o}||={\mathcal{O}}(\epsilon)$. Define the common node set ${\mathcal{N}}_{c}^{1}={\mathcal{N}}_{o}^{1}\cap{\mathcal{N}}_{n}^{1}$, the added node set ${\mathcal{N}}_{a}^{1}={\mathcal{N}}_{n}^{1}\setminus{\mathcal{N}}_{c}^{1}$, and the removed node set ${\mathcal{N}}_{r}^{1}={\mathcal{N}}_{o}^{1}\setminus{\mathcal{N}}_{c}^{1}$. Since $\eta$ is locally diminishing for $v\in{\mathcal{N}}_{a,r}^{1}$, we have a positive value $K_{1}$ such that $\eta={\mathcal{O}}(K_{1}\epsilon)={\mathcal{O}}(\epsilon)$. The pertubation for the assembled matrix ${\bm{M}}$ before/after the particle $p$ crosses an edge is
@@
Similarly, for the perturbation in the assembled vector ${\bm{B}}{\bm{u}}={\bm{P}}{\bm{D}}{\bm{u}}$ before/after the particle crossing is
$||\delta\left({\bm{P}}{\bm{D}}{\bm{u}}\right)||$ $=||\sum_{v\in{\mathcal{N}}_{c}^{1}}\delta(\eta du{\bm{p}})+\sum_{v\in{\mathcal{N}}_{a}^{1}}\eta du{\bm{p}}-\sum_{v\in{\mathcal{N}}_{r}^{1}}\eta du{\bm{p}}||$ (40)
@@
Furthermore, we can establish the following bound for the assembled vector ${\bm{P}}{\bm{D}}{\bm{u}}$:
$||{\bm{P}}{\bm{D}}{\bm{u}}||$ $=||\sum_{v\in{\mathcal{N}}^{1}}\eta du{\bm{p}}||$ (41)
@@
Finally, the perturbation for $\left[\hat{u},\nabla\hat{u}^{T}\right]^{T}$ from (31) is
$\left[\hat{u},\nabla\hat{u}^{T}\right]^{T}$ $=||\delta({\bm{M}}^{-1}{\bm{B}}{\bm{u}})||$ (42)
$=||\delta({\bm{M}}^{-1}{\bm{P}}{\bm{D}}{\bm{u}})||$
$=||\delta{\bm{M}}^{-1}{\bm{P}}{\bm{D}}{\bm{u}}+{\bm{M}}^{-1}\delta\left({\bm{P}}{\bm{D}}{\bm{u}}\right)||$
$\leq||\delta{\bm{M}}^{-1}{\bm{P}}{\bm{D}}{\bm{u}}||+||{\bm{M}}^{-1}\delta\left({\bm{P}}{\bm{D}}{\bm{u}}\right)||$
$\leq||\delta{\bm{M}}^{-1}||\cdot||{\bm{P}}{\bm{D}}{\bm{u}}||+||{\bm{M}}^{-1}||\cdot||\delta\left({\bm{P}}{\bm{D}}{\bm{u}}\right)||$
@@
In the incomplete singular value decomposition of ${\bm{M}}$, the singular values will always be non-negative. And if the surrounding nodes are not degenerate, the minimum singular value $\sigma({\bm{M}})_{\min}$ will always be positive and the condition number of ${\bm{M}}$ is bounded. Therefore, as long as the mesh is of reasonably good quality, both the function value and gradient estimation is ${\mathcal{C}}^{0}$ across the boundary.
Proposition. The function $\eta_{i}$ in (32) is locally diminishing for $\forall i\in{\mathcal{N}}_{a,r}^{1}$.
@@
Formally, we need to prove that for any $i\in{\mathcal{N}}_{a,r}^{1}$, when ${\bm{x}}$ is crossing the edge of a triangle and $||{\bm{x}}^{n}-{\bm{x}}^{o}||={\mathcal{O}}(\epsilon)$, the smoothing function $\eta_{i}={\mathcal{O}}(\epsilon)$.
Denote the edge that the particle is crossing as $e$ and the portion of $||{\bm{x}}^{n}-{\bm{x}}^{o}||$ in the new/old cell as $L^{n,o}$. Trivially,
@@
Then, let the far-away node not on the edge but in the new/old cell be $i_{\text{far}}$ (i.e., $i_{\text{far}}^{n,o}\notin e\,\land\,i_{\text{far}}^{n,o}\in{\mathcal{N}}_{o,n}^{0}$) and the height from a node $i$ to an edge $e$ be $H(i,e)$. Since the height is orthogonal to the edge, we have $H({\bm{x}}^{n,o},e)\leq L^{n,o}={\mathcal{O}}(\epsilon)$. Consider the barycentric coordinate contributed by the far-away node, in the new/old cell respectively, for ${\bm{x}}$:
$B^{n,o}_{i_{\text{far}}}$ $=\frac{H({\bm{x}}^{n,o},e)\cdot||e||}{H(i_{\text{far}}^{n,o},e)\cdot||e||}$ (44)
$=\frac{H({\bm{x}}^{n,o},e)}{H(i_{\text{far}}^{n,o},e)}$
$={\mathcal{O}}(\frac{\epsilon}{H(i_{\text{far}}^{n,o},e)})$
@@
Figure 25: A visual representation of the notations used to prove that $\eta$ diminishes locally, as described in ((44)), for every vertex $i$ within the first ring of neighbors, ${\mathcal{N}}^{1}_{a,r}$. The dashed line denotes the perpendicular height from a given position to the shared edge. Dotted lines are drawn to construct a triangle between the point ${\bm{x}}$ and the edge, facilitating the computation of the barycentric coordinates.
S
Finally, if a node is added/removed during the particle crossing (i.e., $i\in{\mathcal{N}}_{a,r}^{1}$), this means that $i$ is only connected to the far-away nodes $i_{\text{far}}^{n,o}$ but not to the edge $e$; i.e., ${A}_{i,i_{\text{far}}^{n,o}}=1,\forall i\in{\mathcal{N}}^{1}_{a,r}$, otherwise ${A}_{i,n}=0,\forall n\in e\land\forall i\in{\mathcal{N}}^{1}_{a,r}$. Hence,
$\eta$ $=\sum_{n\in{\mathcal{N}}^{0}}B_{n}{A}_{i,n}$ (45)
$=B^{n,o}_{i_{\text{far}}}{A}_{i,i_{\text{far}}^{n,o}}+\sum_{n\in e}B_{n}{A}_{i,n}$
$=B^{n,o}_{i_{\text{far}}}\cdot 1+\sum_{n\in e}B_{n}\cdot 0$
$=B^{n,o}_{i_{\text{far}}}$
$={\mathcal{O}}(\epsilon),\quad\forall i\in{\mathcal{N}}^{1}_{a,r}.$
This concludes the proof.
Appendix C Settings and Analytical Solutions for the Verification Experiments of the Continuous Reconstructions
@@
Appendix D Conservation of the Linear and Affine Momentum When Combined With Affine Particle-in-Cell
@@
$\sum_{i}w_{p,i}^{n}$ $=1,$
$\sum_{i}w_{p,i}^{n}{\bm{x}}_{i}^{n}$ $={\bm{x}}_{p}^{n},$
$\sum_{i}w_{p,i}^{n}({\bm{x}}_{i}^{n}-{\bm{x}}_{p}^{n})$ $=0,$
then the system’s total linear and angular momentum will be conserved when combined with APIC. A simple introduction to APIC is given here for the sake of completeness, while the detailed proof can found in the supplementary document of the original APIC paper [35].
In APIC, mass $m_{p}$, position ${\bm{x}}_{p}$, velocity ${\bm{v}}_{p}$, and an affine matrix ${\bm{B}}_{p}=\sum_{i}w_{p,i}v_{i}({\bm{x}}_{i}-{\bm{x}}_{p})^{T}$ are stored and tracked on particles. Then,
@@
$\begin{split}m_{i}^{n}&=\sum_{p}w_{p,i}^{n}m_{p}\\ {\bm{D}}_{p}^{n}&=\sum_{i}w_{p,i}^{n}({\bm{x}}_{i}^{n}-{\bm{x}}_{p}^{n})({\bm{x}}_{i}^{n}-{\bm{x}}_{p}^{n})^{T}\\ m_{i}^{n}{\bm{v}}_{i}^{n}&=\sum_{p}w_{p,i}^{n}m_{p}({\bm{v}}_{p}^{n}+{\bm{B}}_{p}^{n}({\bm{D}}_{p}^{n})^{-1}({\bm{x}}_{i}^{n}-{\bm{x}}_{p}^{n}))\end{split}$ (48)
@@
$\begin{split}{\bm{v}}_{p}^{n+1}&=\sum_{i}w_{p,i}^{n}\tilde{v}_{i}^{n+1}\\ {\bm{B}}_{p}^{n+1}&=\sum_{i}w_{p,i}^{n}\tilde{v}_{i}^{n+1}({\bm{x}}_{i}^{n}-{\bm{x}}_{p}^{n})^{T},\end{split}$ (49)